# Topological Indices With Degree Sequence $\mathscr{D}$ of Tree

### v1

@@ 3.1. Albertson Index With Degree Sequence $d_{n}\geqslant\dots\geqslant d_{1}$
  • In this case, the sequence is $d=(d_{1},d_{3},d_{4},d_{2},d_{5})$, the constant term is $\lvert d_{3}-d_{4}\rvert+\lvert d_{2}-d_{5}\rvert-2$, then we have,
@@
  • In this case, the sequence is $d=(d_{2},d_{3},d_{1},d_{5},d_{4})$, then we have
@@
  • In this case, the sequence is $d=(d_{2},d_{1},d_{5},d_{3},d_{4})$, then we have
@@
  • In this case, given the sequence as $d=(d_{3},d_{1},d_{2},d_{4},d_{5})$, then we have
@@
  • In this case, given the sequence as $d=(d_{3},d_{2},d_{4},d_{1},d_{5})$, then we have
@@
  • In this case, the sequence is $d=(d_{3},d_{4},d_{1},d_{2},d_{5})$, we notice that in this case the constant term is $|d_{4}-d_{1}|+|d_{2}-d_{5}|-2$, then we have
@@
Actually, respecting the constant terms in each case that shift with the shift of vertices, we notice in Case 1 for $\operatorname{irr}(T_{2}),\operatorname{irr}(T_{3})$, the maximum value is $\operatorname{irr}(T_{3})$ and the minimum value is $\operatorname{irr}(T_{2})$, is available for each case, another aspect to note is that there are repeated results by comparing the vertices with each other, so we restrict ourselves to these comparisons, which they have distributed through three main cases that give us the maximum and minimum values in this sequence, considering the constant terms in each of them.
Thus, we have:
@@
Let be $T$ is a tree of order $n$, degree sequence $d=(d_{1},d_{2},\dots,d_{6})$ where $d_{6}\geqslant\dots\geqslant d_{1}$, then Albertson index over tree $T$ is given by:
@@
Let be $T$ tree of order $n$, a degree sequence is $d=(d_{1},\dots,d_{n})$ where $d_{n}\geqslant\dots\geqslant d_{1}$, then Albertson index of tree $T$ is:
$\operatorname{irr}(T)=d_{1}^{2}+d_{n}^{2}+\sum_{i=2}^{n-1}d_{i}^{2}+\sum_{i=2}^{n-1}d_{i}+d_{n}-d_{1}-2n+2.$
@@
Let be assume the sequence is $d=(d_{1},\dots,d_{n})$ where $d_{n}\geqslant\dots\geqslant d_{1}$, then by considering measures the total absolute difference between consecutive terms are $\sum_{i=1}^{n-1}(d_{i+1}-d_{i})=(d_{2}-d_{1})+(d_{3}-d_{2})+\dots+(d_{n}-d_{n-1})=d_{n}-d_{1}$. In performing the proof of the relationship, we use mathematical induction, where we find that the relationship is clearly true for certain values, so we will prove the validity of the relationship for $n$ and then for $n+1$.
@@
3.2. Sigma Index With Degree Sequence $d_{n}\geqslant\dots\geqslant d_{1}$
@@
Let be $T$ tree of order $n\leqslant 3$, a sequence $d=(d_{1},d_{2},d_{3})$ where $d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index is:
@@
Furthermore, in Lemma 3.8 we provide a sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{4}\geqslant\dots\geqslant d_{1}$ to define Sigma index among tree with determine sequence as we show that in Lemma 3.7 where the vertices is located as random and by respectively the largest and smallest vertex.
@@
Let $T$ be a tree of order $n>0$ and a sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index is:
@@
According to Lemma 3.8,3.7, in Hypothesize 2, we present a preliminaries for the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$ in Lemma LABEL:le.sigma5 Sigma index for determine sequence.
@@
Let $T$ tree of order $n>0$, and let be the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$, then Sigma index of tree is:
$\sigma(T)=\sum_{i=1}^{4}d_{i}(d_{i+1})^{2}+(d_{1}-1)^{3}+(d_{4})^{3}+\sum_{i=1}^{5}(d_{i}-d_{i+1})^{2}.$
@@
According to the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$, we have many cases of position the vertices in tree $T$. In fact, the order in which the elements of this sequence can be realised is 120, and in order to choose the maximum and minimum cases, we get 10 cases, which we can demonstrate as follows:
@@
In this case, we will consider the largest is $d_{1},d_{5}$ by employee $d_{1}$ as constant term, then we have:
$\sigma(T_{1})=\begin{cases}3d_{1}^{3}-2d_{1}^{2}+4d_{1}+d_{2}d_{3}^{2}+d_{4}^{3}+3&d=(d_{1},d_{2},d_{3},d_{4},d_{5}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{2}^{3}+3&d=(d_{1},d_{2},d_{3},d_{5},d_{4}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{3}^{3}+14&d=(d_{1},d_{2},d_{5},d_{3},d_{4}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{3}^{3}+19&d=(d_{1},d_{5},d_{2},d_{3},d_{4}),\\ 3d_{1}^{3}-d_{1}^{2}+4d_{1}+d_{2}d_{3}^{2}+d_{3}d_{4}^{2}+14&d=(d_{1},d_{3},d_{4},d_{5},d_{2}).\end{cases}$
@@
In this case, randomly picking vertex values and considering the elimination of matched results, so we have
$\sigma(T_{2})=\begin{cases}2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{1}d_{2}^{2}+d_{3}^{2}+d_{4}^{3}+10&d=(d_{5},d_{1},d_{2},d_{3},d_{4}),\\ 2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}d_{3}^{2}+d_{2}^{3}+d_{4}^{3}+21&d=(d_{5},d_{1},d_{4},d_{2},d_{3}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{2}d_{3}^{2}+d_{3}d_{4}^{2}+d_{2}^{3}+22&d=(d_{1},d_{5},d_{2},d_{4},d_{3}),\\ 2d_{1}^{3}-4d_{1}^{2}+5d_{1}+d_{1}d_{2}^{2}+d_{2}d_{3}^{2}+d_{4}^{3}+8&d=(d_{5},d_{4},d_{1},d_{3},d_{2}),\\ 2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}d_{3}^{2}+d_{2}^{3}+d_{4}^{3}+24&d=(d_{5},d_{2},d_{4},d_{1},d_{3}).\end{cases}$
@@
In this case, To analyse these cases, we will debate each case with the others and then extract the maximum case value and demonstrate it through the following discussion as:
$\sigma(T_{3})=\begin{cases}d_{1}^{3}+2d_{1}^{2}+d_{1}+d_{2}d_{3}^{2}-d_{1}d_{2}^{2}+d_{4}^{3}-d_{3}d_{4}^{2}-d_{2}^{3},&d_{1}\geqslant 2\\ d_{1}^{3}-3d_{1}^{2}-d_{1}+d_{1}d_{2}^{2}-d_{2}d_{3}^{2}+d_{2}^{3}-11,&d_{1}\geqslant 3\\ d_{1}^{2}-2d_{1}+d_{3}d_{4}^{2}-d_{3}^{2}+d_{3}^{3}-d_{4}^{3}+4&d_{1}\geqslant 3\\ d_{1}^{2}-2d_{1}+d_{3}d_{4}^{2}-d_{3}^{2}+d_{3}^{3}-d_{4}^{3}+9&d_{1}\geqslant 2\\ d_{1}^{3}+4d_{1}^{2}-d_{1}+d_{2}d_{3}^{2}-d_{1}d_{2}^{2}+d_{3}d_{4}^{2}-d_{3}^{2}-d_{4}^{3}+4&d_{1}\geqslant 3\\ d_{1}d_{2}^{2}-d_{2}d_{3}^{2}+d_{3}^{2}-d_{2}^{3}-14&d_{1}\geqslant 3\\ -d_{1}^{2}+2d_{1}-d_{3}d_{4}^{2}+d_{4}^{3}-1&d_{1}\geqslant 5\\ -2d_{1}+d_{3}d_{4}^{2}-d_{1}d_{2}^{2}+d_{2}^{3}-d_{4}^{3}+14&d_{1}\geqslant 3.\end{cases}$
@@
$\sigma(T)=\begin{cases}\sigma_{\max}(T)=2d_{1}^{3}-4d_{1}^{2}+5d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{4}^{3}+8\\ \sigma_{\min}(T)=2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{2}d_{3}^{2}+d_{3}^{3}+19.\end{cases}$
As desire. ∎ In Table 1 we show some values of degree sequence according to Hypothesize 2 for analyse these cases with updated calculations the modified formulas. Each row pairs vertices differing only in $d_{5}$ as we show that.
Table 1. Values of $\sigma_{\max}(T)$ and $\sigma_{\min}(T)$ for Selected Parameters
$(d_{1},\ldots,d_{5})$ $\sigma_{\max}(T)$ $\sigma_{\min}(T)$ $(d_{1},\ldots,d_{5})$ $\sigma_{\max}(T)$ $\sigma_{\min}(T)$
(4,7,12,15,20) 6363 3027 (4,7,12,15,23) 6363 3027
(4,7,12,17,20) 8669 3027 (4,7,12,17,23) 8669 3027
(4,7,14,15,20) 7887 5955 (4,7,14,15,23) 7887 5955
(4,10,12,17,20) 8873 3663 (4,10,12,17,23) 8873 3663
(4,10,14,15,20) 8091 6591 (4,10,14,15,23) 8091 6591
(4,10,14,17,20) 10397 6591 (4,10,14,17,23) 10397 6591
(6,7,12,15,20) 9885 5629 (6,7,12,15,23) 9885 5629
(6,7,12,17,20) 13191 5629 (6,7,12,17,23) 13191 5629
(6,7,14,15,20) 11409 8557 (6,7,14,15,23) 11409 8557
(6,7,14,17,20) 14715 8557 (6,7,14,17,23) 14715 8557
(6,10,14,15,20) 11613 9193 (6,10,14,15,23) 11613 9193
(6,10,14,17,20) 14919 9193 (6,10,14,17,23) 14919 9193
@@
Let be a sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index is:
@@
We will discuss two basic conditions in the proof, one of which leads to the other, as it is fulfilled under certain conditions as follows:
@@
In this case we assume the sequence is $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, in this case the term is $d_{i}=d_{i-1}+1$, the number of possible cases in this case is 720 cases, some of which give similar results and some of which do not fulfil the condition and give negative values, so we took 10 cases and compared them with the rest of the cases and reviewed the results of this comparison through cases from case 1 to case 10.
@@
In this case we assume the sequence is $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, this case is a generalisation of the previous case but with no specific conditions, just a normal order.
@@
in this case we have
@@
in this case we have the constant term is differ $d_{1}$, by considering $\beta=d_{2}d_{3}^{2}+d_{3}^{3}+2d_{3}^{2}+d_{3}+d_{5}^{3}$, $\gamma=d_{4}d_{5}^{2}+2d_{4}^{2}+d_{2}d_{4}^{2}$ and $\zeta=d_{1}d_{2}^{2}+d_{3}d_{5}^{2}+d_{5}^{2}$ then we have:
@@
Now, let be considered the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$, then we have according Case b where considering $k>0,k\in\mathbb{N}$,and the number of pendent vertices $n\geq 1$ then:
@@
In this case, we have the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$, then we have:
@@
In this case, we have the sequence $d=(d_{6},d_{4},d_{2},d_{1},d_{3},d_{5})$, then:
@@
Let be $T$ a tree of order $n$ define a sequence $d=(d_{1},\dots,d_{n})$ where $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{n}$, then Albertson index among tree $T$ is:
@@
Let $T$ be a tree of order $n>0$, define degree sequence $d=(d_{1},\dots,d_{n})$ where $d_{n}\geqslant\dots\geqslant d_{1}$, then Sigma index among tree given as:

### v2

@@ 3.1. Albertson Index With Non- Increasing Degree Sequence
  • In this case, the sequence is $\mathscr{D}=(d_{1},d_{3},d_{4},d_{2},d_{5})$, the constant term is $\lvert d_{3}-d_{4}\rvert+\lvert d_{2}-d_{5}\rvert-2$, then we have,
@@
  • In this case, the sequence is $\mathscr{D}=(d_{2},d_{3},d_{1},d_{5},d_{4})$, then we have
@@
  • In this case, the sequence is $\mathscr{D}=(d_{2},d_{1},d_{5},d_{3},d_{4})$, then we have
@@
  • In this case, given the sequence as $\mathscr{D}=(d_{3},d_{1},d_{2},d_{4},d_{5})$, then we have
@@
  • In this case, given the sequence as $\mathscr{D}=(d_{3},d_{2},d_{4},d_{1},d_{5})$, then we have
@@
  • In this case, the sequence is $\mathscr{D}=(d_{3},d_{4},d_{1},d_{2},d_{5})$, we notice that in this case the constant term is $|d_{4}-d_{1}|+|d_{2}-d_{5}|-2$, then we have
@@
Actually, respecting the constant terms in each case that shift with the shift of vertices, we notice in Case 1 for $\operatorname{irr}(T_{2}),\operatorname{irr}(T_{3})$, the maximum value is $\operatorname{irr}(T_{3})$ and the minimum value is $\operatorname{irr}(T_{2})$, is available for each case, another aspect to note is that there are repeated results by comparing the vertices with each other, so we restrict ourselves to these comparisons, which they have distributed through three main cases that give us the maximum and minimum values in this sequence, considering the constant terms in each of them. Thus, we have:
@@
Let $T$ be is a tree of order $n$, let $\mathscr{D}=(d_{1},\dots,d_{6})$ be a non-decreasing degree sequence where $d_{6}\geqslant\dots\geqslant d_{1}$, then Albertson index over tree $T$ is given by:
@@
Let $T$ be tree of order $n$, let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a non-decreasing degree sequence where $d_{n}\geqslant\dots\geqslant d_{1}$, then Albertson index of tree $T$ is:
$\operatorname{irr}(T)=d_{1}^{2}+d_{n}^{2}+\sum_{i=2}^{n-1}d_{i}^{2}+\sum_{i=2}^{n-1}d_{i}+d_{n}-d_{1}-2n-2.$
@@
Let be assume the sequence is $\mathscr{D}=(d_{1},\dots,d_{n})$ where $d_{n}\geqslant\dots\geqslant d_{1}$, then by considering measures the total absolute difference between consecutive terms are $\sum_{i=1}^{n-1}(d_{i+1}-d_{i})=(d_{2}-d_{1})+(d_{3}-d_{2})+\dots+(d_{n}-d_{n-1})=d_{n}-d_{1}$. In performing the proof of the relationship, we use mathematical induction, where we find that the relationship is clearly true for certain values, so we will prove the validity of the relationship for $n$ and then for $n+1$.
@@
3.2. Sigma Index With Non-Decreasing Degree Sequence
@@
Let $T$ be tree of order $n\leqslant 3$, let $\mathscr{D}=(d_{1},d_{2},d_{3})$ be a degree sequence where $d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index is:
@@
Furthermore, in Lemma 3.8 we provide $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4})$ a degree sequence where $d_{4}\geqslant\dots\geqslant d_{1}$ to define Sigma index among tree with determine sequence as we show that in Lemma 3.7 where the vertices is located as random and by respectively the largest and smallest vertex.
@@
Let $T$ be a tree of order $n>0$ and let $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4})$ be a degree sequence where $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index is:
@@
According to Lemma 3.8,3.7, in Hypothesize 2, we present a preliminaries for the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$ as we show that.
@@
Let $T$ be a tree of order $n>0$, and let $\mathscr{D}=(d_{1},\dots,d_{5})$ be a degree sequence where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$, then Sigma index of $T$ given by:
$\sigma(T)=\sum_{i=1}^{3}d_{i}(d_{i+1})^{2}+(d_{1}-1)^{3}+(d_{4})^{3}+\sum_{i=1}^{4}(d_{i}-d_{i+1})^{2}.$
@@
According to the sequence $\mathscr{D}=(d_{1},\dots,d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$, we have many cases of position the vertices in tree $T$. In fact, the order in which the elements of this sequence can be realised is 120, and in order to choose the maximum and minimum cases, we get 10 cases, which we can demonstrate as follows:
@@
in this case, we have the sequence is $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4},d_{5})$, then we have:
$\sigma(T)$ $=(d_{1}-1)^{3}+(d_{1}-1)(d_{1})^{2}+d_{1}(d_{1}+1)^{2}+d_{2}(d_{3})^{2}+(d_{4})^{3}+4$
$=3d_{1}^{3}-2d_{1}^{2}+4d_{1}+d_{2}d_{3}^{2}+d_{4}^{3}+3.$
@@
in this case, we have the sequence is $\mathscr{D}=(d_{1},d_{2},d_{3},d_{5},d_{4})$, then we have:
$\sigma(T)$ $=(d_{1}-1)^{3}+(d_{1}-1)(d_{1})^{2}+d_{1}(d_{2})^{2}+d_{3}(d_{4})^{2}+(d_{2})^{3}+4$
$=2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{2}^{3}+3.$
@@
in this case, we have the sequence is $\mathscr{D}=(d_{1},d_{2},d_{5},d_{3},d_{4})$, then we have:
$\sigma(T)$ $=(d_{1}-1)^{3}+(d_{1}-1)(d_{1})^{2}+d_{3}(d_{4})^{2}+d_{1}(d_{2})^{2}+(d_{3})^{3}+15$
$=2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{3}^{3}+14.$
Case 4:
in this case, we have the sequence is $\mathscr{D}=(d_{1},d_{5},d_{2},d_{3},d_{4})$, then we have:
$\sigma(T)$ $=(d_{1}-1)^{3}+(d_{3})(d_{4})^{2}+(d_{1}-1)(d_{1})^{2}+d_{1}(d_{2})^{2}+(d_{3})^{3}+20$
$=2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{3}^{3}+19.$
Case 5:
in this case, we have the sequence is $\mathscr{D}=(d_{1},d_{3},d_{4},d_{5},d_{2})$, then we have:
$\sigma(T)$ $=(d_{1}-1)^{3}+d_{1}(d_{1}+1)^{2}+d_{2}(d_{3})^{2}+d_{3}(d_{4})^{2}+(d_{1})^{3}+15$
$=3d_{1}^{3}-d_{1}^{2}+4d_{1}+d_{2}d_{3}^{2}+d_{3}d_{4}^{2}+14.$
Case 6:
in this case, we have the sequence is $\mathscr{D}=(d_{5},d_{1},d_{2},d_{3},d_{4})$, then we have:
$\sigma(T)$ $=(d_{4})^{3}+(d_{1}-2)(d_{1}-1)^{2}+(d_{1}-1)(d_{1})^{2}+d_{1}(d_{2})^{2}+(d_{3})^{2}+12$
$=2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{1}d_{2}^{2}+d_{3}^{2}+d_{4}^{3}+10.$
Case 7:
in this case, we have the sequence is $\mathscr{D}=(d_{5},d_{1},d_{4},d_{2},d_{3})$, then we have:
$\sigma(T)$ $=(d_{4})^{3}+(d_{1}-2)(d_{1}-1)^{2}+(d_{2})(d_{3})^{2}+(d_{1}-1)(d_{1})^{2}+(d_{2})^{3}+23$
$=2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}d_{3}^{2}+d_{2}^{3}+d_{4}^{3}+21.$
Case 8:
in this case, we have the sequence is $\mathscr{D}=(d_{1},d_{5},d_{2},d_{4},d_{3})$, then we have:
$\sigma(T)$ $=(d_{1}-1)^{3}+(d_{3})(d_{4})^{2}+(d_{1}-1)(d_{1})^{2}+(d_{2})(d_{3})^{2}+(d_{2})^{3}+23$
$=2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{2}d_{3}^{2}+d_{3}d_{4}^{2}+d_{2}^{3}+22.$
Case 9:
in this case, we have the sequence is $\mathscr{D}=(d_{5},d_{4},d_{1},d_{3},d_{2})$, then we have:
$\sigma(T)$ $=(d_{4})^{3}+(d_{2})(d_{3})^{2}+(d_{1}-2)(d_{1}-1)^{2}+(d_{1})(d_{2})^{2}+(d_{1})^{3}+10$
$=2d_{1}^{3}-4d_{1}^{2}+5d_{1}+d_{1}d_{2}^{2}+d_{2}d_{3}^{2}+d_{4}^{3}+8.$
Case 10:
in this case, we have the sequence is $\mathscr{D}=(d_{5},d_{2},d_{4},d_{1},d_{3})$, then we have:
$\sigma(T)$ $=(d_{4})^{3}+(d_{1}-1)(d_{1})^{2}+(d_{2})(d_{3})^{2}+(d_{1}-2)(d_{1}-1)^{2}+(d_{2})^{3}+26$
$=2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}d_{3}^{2}+d_{2}^{3}+d_{4}^{3}+24.$
To analyse these cases, we will debate each case with the others and then extract the maximum case value and demonstrate it through the following discussion as:
  • From Case 1, we have $\sigma(T)=d_{1}^{3}+2d_{1}^{2}+d_{1}+d_{2}d_{3}^{2}-d_{1}d_{2}^{2}+d_{4}^{3}-d_{3}d_{4}^{2}-d_{2}^{3}$,
  • From Case 2, we have $\sigma(T)=-d_{1}^{3}-3d_{1}^{2}-d_{1}+d_{1}d_{2}^{2}-d_{2}d_{3}^{2}+d_{2}^{3}-11$ if and only if $d_{1}\geqslant 3$,
  • From Case 3, we have $\sigma(T)=d_{1}^{2}-2d_{1}+d_{3}d_{4}^{2}-d_{3}^{2}+d_{3}^{3}-d_{4}^{3}+4$,
  • From Case 4, we have $\sigma(T)=d_{1}^{2}-2d_{1}+d_{3}d_{4}^{2}-d_{3}^{2}+d_{3}^{3}-d_{4}^{3}+9$,
  • From Case 5, we have $\sigma(T)=d_{1}^{3}+4d_{1}^{2}-d_{1}+d_{2}d_{3}^{2}-d_{1}d_{2}^{2}+d_{3}d_{4}^{2}-d_{3}^{2}-d_{4}^{3}+4$,
  • from Case 6, we have $\sigma(T)=d_{1}d_{2}^{2}-d_{2}d_{3}^{2}+d_{3}^{2}-d_{2}^{3}-14$,
  • From Case 7, we have $\sigma(T)=-d_{1}^{2}+2d_{1}-d_{3}d_{4}^{2}+d_{4}^{3}-1$,
  • From Case 8,9,10 we have $\sigma(T)=-2d_{1}+d_{3}d_{4}^{2}-d_{1}d_{2}^{2}+d_{2}^{3}-d_{4}^{3}+14$.
@@
$\sigma(T)=\begin{cases}\sigma_{\max}(T)=d_{1}^{3}+2d_{1}^{2}+d_{1}+d_{2}d_{3}^{2}-d_{1}d_{2}^{2}+d_{4}^{3}-d_{3}d_{4}^{2}-d_{2}^{3}\\ \sigma_{\min}(T)=-d_{1}^{2}+2d_{1}-d_{3}d_{4}^{2}+d_{4}^{3}-1.\end{cases}$
As desire. ∎
@@
let $\mathscr{D}=(d_{1},\dots,d_{5})$ be a degree sequence where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index of $T$ is:
@@
We will discuss two basic conditions in the proof, one of which leads to the other, as it is fulfilled under certain conditions according to Hypothesize 2 and Lemma 3.8 as follows:
@@
In this case, we assume the sequence is $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, in this case the term is $d_{i}=d_{i-1}+1$, the number of possible cases in this case is 720 cases, some of which give similar results and some of which do not fulfil the condition and give negative values, so we took 10 cases and compared them with the rest of the cases and reviewed the results of this comparison through cases from case 1 to case 10.
@@
In this case, we assume the sequence is $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, this case is a generalisation of the previous case but with no specific conditions, just a normal order.
@@
In this case, we have
@@
In this case, we have the constant term is differ $d_{1}$, by considering $\beta=d_{2}d_{3}^{2}+d_{3}^{3}+2d_{3}^{2}+d_{3}+d_{5}^{3}$, $\gamma=d_{4}d_{5}^{2}+2d_{4}^{2}+d_{2}d_{4}^{2}$ and $\zeta=d_{1}d_{2}^{2}+d_{3}d_{5}^{2}+d_{5}^{2}$ then we have:
@@
Now, let be considered the sequence $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$, then we have according Case b where considering $k>0,k\in\mathbb{N}$,and the number of pendent vertices $n\geq 1$ then:
@@
In this case, we have the sequence $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$, then we have:
@@
In this case, we have the sequence $\mathscr{D}=(d_{6},d_{4},d_{2},d_{1},d_{3},d_{5})$, then:
@@
Let $T$ be a tree of order $n$, let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence where $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{n}$, then Albertson index among tree $T$ is:
@@
Let $T$ be a tree of order $n>0$, let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence where $d_{n}\geqslant\dots\geqslant d_{1}$, then Sigma index among tree given as: